# Observing network dynamics through sentinel nodes

### v1

@@ Acknowledgements
Optimized -- Random $-1.748$ $[-1.794,-1.701]$ $<10^{-7}$
Weight-optimized -- Random $-1.728$ $[-1.775,-1.682]$ $<10^{-7}$
Weight-optimized -- Optimized $0.019$ $[-0.027,0.066]$ $0.591$
Appendix S9 Sentinel node set optimization on weighted networks
We collected ten weighted networks from online repositories. From the KONECT project [61], we used a dominance network of Japanese macaques [76] ($N=62$, $M=1167$), a friendship network of high school students in Illinois, USA [77] ($N=70$, $M=274$), a friendship network of Australian university students [78] ($N=217$, $M=1839$), and the proximity network from the main text but with edge weights retained [62] ($N=410$, $M=2765$). From the Netzschleuder repository [79], we used a network of interpersonal contacts between windsurfers [80] ($N=43$, $M=336$), a network of contacts between individuals involved in the train bombing in 2004 in Madrid, Spain [81] ($N=64$, $M=243$), a network of drug interactions gathered from health records in Blumenau, Brazil [82] ($N=75$, $M=181$), a coauthorship network [83] ($N=379$, $M=914$)), a neuronal network of Caenorhabditis elegans [84] ($N=460$, $M=1432$), and a product export network [85] ($N=774$, $M=1779$). As we did for unweighted networks in the main text, we coerced each network to be undirected, discarded any self-edges, and analyzed only the largest connected component.
Figure S15: Approximation error on weighted networks for the coupled double-well dynamics for node sets with $n=\lfloor\ln N\rfloor$ nodes. We show the results for optimized (blue circles), degree-preserving (orange triangles), and completely random (green squares) node sets on all networks. Each marker represents the approximation error for one node set. The different panels show the results for different networks. (a) Windsurfer contact (with $N=43$ nodes). (b) Macaque dominance ($N=62$). (c) Terrorist contact ($N=64$). (d) High school friendship ($N=70$). (e) Drug interaction ($N=75$). (f) University friendship ($N=217$). (g) Coauthorship ($N=379$). (h) Weighted proximity ($N=410$). (i) Neuronal ($N=460$). (j) Export ($N=774$).
To systematically verify that the sentinel node approximation performs better than the other node set selection methods, we ran the same statistical analysis as the one we used for unweighted networks (see section S4). The ANOVA results were similar to those for unweighted networks (model $R^{2}=0.73$; dynamics: $df=3$, $F=10300$, $p<10^{-7}$; network: $df=9$, $F=966.3$, $p<10^{-7}$; node set type: $df=6$, $F=6243$, $p<10^{-7}$). The optimized node sets had an average approximation error that was 960 times smaller than completely random node sets ($b=-6.867$, $e^{b}=0.001$, $p<10^{-7}$) and 262 times smaller than degree-preserving node sets (difference in coefficients: $-5.569$, $e^{b}=0.004$, $p<10^{-7}$); see Table S6. We note that the ordering of node set types in terms of approximation error is similar to that for unweighted networks, except that there is less difference between the heuristic node sets and the completely random node sets in the case of weighted networks. For example, both the $k$-constrained and community-based node sets are not significantly different from the completely random node sets.
Difference CI $p$
Optimized -- Random $-6.867$ $[-6.999,-6.735]$ $<10^{-7}$
Degree-preserving -- Random $-1.298$ $[-1.430,-1.166]$ $<10^{-7}$
$k$-constrained -- Random $-0.093$ $[-0.225,0.039]$ $0.370$
$k$-quantiled -- Random $-0.407$ $[-0.539,-0.275]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Random $-0.196$ $[-0.328,-0.064]$ $2.36\times 10^{-4}$
Community-based -- Random $-0.100$ $[-0.232,0.032]$ $0.283$
Degree-preserving -- Optimized $5.569$ $[5.437,5.701]$ $<10^{-7}$
$k$-constrained -- Optimized $6.774$ $[6.642,6.906]$ $<10^{-7}$
$k$-quantiled -- Optimized $6.460$ $[6.328,6.592]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Optimized $6.670$ $[6.538,6.802]$ $<10^{-7}$
Community-based -- Optimized $6.767$ $[6.635,6.899]$ $<10^{-7}$
$k$-constrained -- Degree-preserving $1.205$ $[1.073,1.337]$ $<10^{-7}$
$k$-quantiled -- Degree-preserving $0.891$ $[0.759,1.023]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Degree-preserving $1.101$ $[0.969,1.233]$ $<10^{-7}$
Community-based -- Degree-preserving $1.198$ $[1.066,1.330]$ $<10^{-7}$
$k$-quantiled -- $k$-constrained $-0.314$ $[-0.446,-0.182]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- $k$-constrained $-0.104$ $[-0.236,0.028]$ $0.238$
Community-based -- $k$-constrained $-0.007$ $[-0.139,0.125]$ $>0.999$
$k_{\rm nn}$-constrained -- $k$-quantiled $0.210$ $[0.078,0.343]$ $5.41\times 10^{-5}$
Community-based -- $k$-quantiled $0.307$ $[0.175,0.439]$ $<10^{-7}$
Community-based -- $k_{\rm nn}$-constrained $0.097$ $[-0.035,0.229]$ $0.318$
The performance of our algorithm on weighted networks is also similar to that on unweighted networks when we do not know the test dynamics. In fact, by running the same analysis as that for the unweighted networks (see section S7), we find similar ANOVA results ($R^{2}=0.68$; training dynamics: $df=3$, $F=4440$, $p<10^{-7}$; test dynamics: $df=3$, $F=38634$, $p<10^{-7}$; network: $df=9$, $F=3490$, $p<10^{-7}$; node set type: $df=6$, $F=2391$, $p<10^{-7}$). The average approximation error for the optimized node sets is 9.67 times smaller than for completely random node sets ($b=-2.269$, $e^{b}=0.096$, $p<10^{-7}$; Table S7) and 3.66 times smaller than degree-preserving node sets (difference in coefficients: $-1.297$, $e^{b}=0.273$, $p<10^{-7}$).
Difference CI $p$
Optimized -- Random $-2.269$ $[-2.339,-2.199]$ $<10^{-7}$
Degree-preserving -- Random $-0.972$ $[-1.042,-0.903]$ $<10^{-7}$
$k$-constrained -- Random $-0.093$ $[-0.162,-0.023]$ $1.63\times 10^{-3}$
$k$-quantiled -- Random $-0.407$ $[-0.476,-0.337]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Random $-0.196$ $[-0.266,-0.127]$ $<10^{-7}$
Community-based -- Random $-0.100$ $[-0.169,-0.030]$ $4.79\times 10^{-4}$
Degree-preserving -- Optimized $1.297$ $[1.227,1.366]$ $<10^{-7}$
$k$-constrained -- Optimized $2.176$ $[2.107,2.246]$ $<10^{-7}$
$k$-quantiled -- Optimized $1.862$ $[1.793,1.932]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Optimized $2.073$ $[2.003,2.142]$ $<10^{-7}$
Community-based -- Optimized $2.169$ $[2.100,2.239]$ $<10^{-7}$
$k$-constrained -- Degree-preserving $0.879$ $[0.810,0.949]$ $<10^{-7}$
$k$-quantiled -- Degree-preserving $0.565$ $[0.496,0.635]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Degree-preserving $0.776$ $[0.706,0.845]$ $<10^{-7}$
Community-based -- Degree-preserving $0.873$ $[0.803,0.942]$ $<10^{-7}$
$k$-quantiled -- $k$-constrained $-0.314$ $[-0.384,-0.245]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- $k$-constrained $-0.104$ $[-0.173,-0.034]$ $2.24\times 10^{-4}$
Community-based -- $k$-constrained $-0.007$ $[-0.076,0.063]$ $>0.999$
$k_{\rm nn}$-constrained -- $k$-quantiled $0.210$ $[0.141,0.280]$ $<10^{-7}$
Community-based -- $k$-quantiled $0.307$ $[0.238,0.377]$ $<10^{-7}$
Community-based -- $k_{\rm nn}$-constrained $0.097$ $[0.027,0.166]$ $8.04\times 10^{-4}$
Appendix S10 Sentinel node set optimization on directed networks
We collected ten directed networks from the Netzschleuder repository [79]: a freshwater trophic network [86] ($N=109$, $M=717$), a social network of physicians [87] ($N=117$, $M=542$), an email network from a manufacturing company [88] ($N=167$, $M=5783$), a dependency network of the Flamingo software [89] ($N=228$, $M=497$), a transcription network of the bacterium Escherichia coli [90] ($N=328$, $M=456$), a transcription network of the yeast Saccharomyces cerevisiae [91] ($N=664$, $M=1066$), a network of US air carrier flights in 2020 [92] ($N=806$, $M=11924$), a dependency network of the JUNG software [89] ($N=879$, $M=2051$), the directed version of the university email network from the main text [65] ($N=1133$, $M=10902$), and a network of air routes preferred by the US Federal Aviation Administration [93] ($N=1226$, $M=2613$). We discarded any self-loops and the multiplicity of any multi-edges.
Figure S16: Approximation error for the coupled double-well dynamics on directed networks with $n=\lfloor\ln N\rfloor$ nodes. (a) Trophic. (b) Physicians. (c) Email (manufacturer). (d) Flamingo. (e) Escherichia coli. (f) Saccharomyces cerevisiae. (g) Flights. (h) JUNG. (i) Email (university). (j) Air route.
We verified our observations with an ANOVA (model $R^{2}=0.60$; dynamics: $df=3$, $F=5048$, $p<10^{-7}$; network: $df=9$, $F=1925$, $p<10^{-7}$; node set type: $df=6$, $F=1541$, $p<10^{-7}$). On average, the approximation error for optimized node sets was 211 times smaller than completely random node sets ($b=-5.351$, $e^{b}=0.005$, $p<10^{-7}$) and 62.4 times smaller than degree-preserving node sets (difference in coefficients: $-4.134$, $e^{b}=0.016$, $p<10^{-7}$); see Table S8.
Difference CI $p$
Optimized -- Random $-5.351$ $[-5.563,-5.139]$ $<10^{-7}$
Degree-preserving -- Random $-1.217$ $[-1.429,-1.005]$ $<10^{-7}$
$k$-constrained -- Random $-0.081$ $[-0.293,0.131]$ 0.921
$k$-quantiled -- Random $-0.362$ $[-0.574,-0.150]$ $1.00\times 10^{-5}$
$k_{\rm nn}$-constrained -- Random $-0.045$ $[-0.257,0.167]$ 0.996
Community-based -- Random $0.347$ $[0.134,0.561]$ $3.19\times 10^{-5}$
Degree-preserving -- Optimized $4.134$ $[3.922,4.346]$ $<10^{-7}$
$k$-constrained -- Optimized $5.270$ $[5.058,5.482]$ $<10^{-7}$
$k$-quantiled -- Optimized $4.989$ $[4.777,5.201]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Optimized $5.306$ $[5.094,5.518]$ $<10^{-7}$
Community-based -- Optimized $5.699$ $[5.486,5.912]$ $<10^{-7}$
$k$-constrained -- Degree-preserving $1.136$ $[0.924,1.348]$ $<10^{-7}$
$k$-quantiled -- Degree-preserving $0.855$ $[0.643,1.067]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Degree-preserving $1.172$ $[0.960,1.384]$ $<10^{-7}$
Community-based -- Degree-preserving $1.564$ $[1.351,1.778]$ $<10^{-7}$
$k$-quantiled -- $k$-constrained $-0.281$ $[-0.493,-0.069]$ $1.80\times 10^{-3}$
$k_{\rm nn}$-constrained -- $k$-constrained $0.036$$[0.176,0.248]$ 0.999
Community-based -- $k$-constrained $0.428$ $[0.215,0.641]$ $10^{-7}$
$k_{\rm nn}$-constrained -- $k$-quantiled $0.317$ $[0.105,0.529]$ $2.15\times 10^{-4}$
Community-based -- $k$-quantiled $0.709$ $[0.496,0.922]$ $<10^{-7}$
Community-based -- $k_{\rm nn}$-constrained $0.393$ $[0.179,0.606]$ $1.2\times 10^{-6}$
The ANOVA for directed networks in the case when we do not know the test dynamics fits the data reasonably well (model $R^{2}=0.53$; training dynamics: $df=3$, $F=1384$, $p<10^{-7}$; test dynamics: $df=3$, $F=10888$, $p<10^{-7}$; network: $df=9$, $F=6397$, $p<10^{-7}$; node set type: $df=6$, $F=261$, $p<10^{-7}$). As is the case for undirected networks, the performance of optimized node sets is worse when the test dynamics is different from the training dynamics, but optimized node sets still outperform all other node sets. Optimized node sets had on average 3.02 times lower approximation error than completely random node sets ($b=-1.104$, $e^{b}=0.332$, $p<10^{-7}$) and 1.81 times smaller error than degree-preserving node sets (difference in coefficients: $-0.595$, $e^{b}=0.552$, $p<10^{-7}$); see Table S9.
Difference CI $p$
Optimized -- Random $-1.104$ $[-1.224,-0.984]$ $<10^{-7}$
Degree-preserving -- Random $-0.508$ $[-0.628,-0.388]$ $<10^{-7}$
$k$-constrained -- Random $-0.081$ $[-0.201,0.039]$ 0.422
$k$-quantiled -- Random $-0.362$ $[-0.482,-0.242]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Random $-0.045$ $[-0.165,0.075]$ 0.925
Community-based -- Random $0.351$ $[0.231,0.472]$ $<10^{-7}$
Degree-preserving -- Optimized $0.595$ $[0.475,0.715]$ $<10^{-7}$
$k$-constrained -- Optimized $1.023$ $[0.903,1.143]$ $<10^{-7}$
$k$-quantiled -- Optimized $0.742$ $[0.622,0.862]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Optimized $1.058$ $[0.939,1.178]$ $<10^{-7}$
Community-based -- Optimized $1.455$ $[1.334,1.575]$ $<10^{-7}$
$k$-constrained -- Degree-preserving $0.427$ $[0.308,0.547]$ $<10^{-7}$
$k$-quantiled -- Degree-preserving $0.146$ $[0.027,0.266]$ $5.86\times 10^{-3}$
$k_{\rm nn}$-constrained -- Degree-preserving $0.463$ $[0.343,0.583]$ $<10^{-7}$
Community-based -- Degree-preserving $0.859$ $[0.739,0.980]$ $<10^{-7}$
$k$-quantiled -- $k$-constrained $-0.281$ $[-0.401,-0.161]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- $k$-constrained $0.036$ $[-0.084,0.156]$ 0.976
Community-based -- $k$-constrained $0.432$ $[0.311,0.552]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- $k$-quantiled $0.317$ $[0.197,0.437]$ $<10^{-7}$
Community-based -- $k$-quantiled $0.713$ $[0.592,0.834]$ $<10^{-7}$
Community-based -- $k_{\rm nn}$-constrained $0.396$ $[0.276,0.517]$ $<10^{-7}$

### v3

@@ C Transfer learning
Optimized - Random 1.748-1.748 [1.794,1.701][-1.794,-1.701] $10^{-7}$
Weight-optimized - Random 1.728-1.728 [1.775,1.682][-1.775,-1.682] $10^{-7}$
Weight-optimized - Optimized $0.019$ [0.027,0.066][-0.027,0.066] $0.591$
Appendix S17 Time complexity
A Node set optimization algorithm
In the simulated annealing algorithm to determine the sentinel node set, $S$, we iterate the steps of the algorithm $h_{\max}=50N$ times. Each step of the algorithm consists in selecting a node to be kicked out from $S$ and a different node to be recruited to $S$, both of which need $O(\ln N)$ time by a standard binary search algorithm. Each step also involves the computation of $\varepsilon$ for the tentative new sentinel node set, which requires $O(N)$ time. Therefore, we expect that the time complexity of the entire optimization algorithm is $O(N^{2})$.
We investigated the actual runtime of the optimization algorithm for all the combinations of dynamics and network. We show the results in Fig. S32. Each symbol in the figure represents the largest runtime among 100 independent runs of the algorithm. The figure indicates that the runtime is $O(N^{2})$ or somewhat smaller.
In contrast, the GBB reduction runs in $O(N)$ time, which is devoted to the calculation of the first and second moment of the node’s degree, yielding $\beta_{\text{eff}}$. The DART requires the leading eigenvalue and eigenvector, which one can obtain by the power method. One iteration of the power method consists of the multiplication of the adjacency matrix and a vector of size $N$, which costs $O(N^{2})$ and $O(N)$ time for dense and sparse networks, respectively, followed by normalization. This multiplication sets the time complexity of DART.
Figure S32: Runtime of the simulated annealing algorithm to optimize the sentinel node set for each of the four dynamics models and each of the 20 networks. A symbol represents the largest runtime among 100 runs for a pair of the dynamics model and network. We used Intel Xeon Gold 6448Y processors with 32 central processing units, 64 threads, and 512 GB of memory.
For two arbitrarily selected networks, we also investigated the runtime as a function of $n$, the number of nodes in the sentinel node set. We show in Fig. S33a the runtime of the simulated annealing algorithm for the Proximity network. Each circle represents the runtime for a single run of the simulated annealing. We used the coupled double-well dynamics. We find that the runtime only modestly increases as $n$ increases. The number of iterations of the algorithm, $h_{\max}$, necessary for sufficient convergence of the approximation error, $\varepsilon$, may depend on $n$. Therefore, we also examined $\varepsilon$ as a function of $h_{\max}$ in Fig. S33b for some representative values of $n$. We recall that we set $h_{\max}=50N$ in generating Fig. S33a as well as all the other results in the main text and the SI. The vertical solid line in Fig. S33b shows $h_{\max}=50N$. The figure suggests that, for all the values of $n$ shown in the figure, $h_{\max}=50N$ is large enough for the convergence of the simulated annealing algorithm. Based on Fig. S33a and b, we conclude that our optimization algorithm is not excessively slow as $n$ increases. We have verified that these results remain qualitatively the same for another larger network (see Fig. S33c and d).
Figure S33: Runtime of the simulated annealing algorithm as a function of the number of nodes in the sentinel node set, $n$. (a) Runtime for the Proximity network. A symbol represents one run of the simulated annealing algorithm. We carried out five runs for each $n$ value used in the figure. (b) Approximation error, $\varepsilon$, as a function of $h_{\max}$ for the Proximity network. (c) Runtime for the Email network. (d) Approximation error for the Email network. We used the coupled double-well dynamics. We set $h_{\max}=50N$ in (a) and (c). See the caption of Fig. S32 for the hardware used.
B Dynamics simulations
For practical application of transfer learning, one needs to select a dynamics and numerically simulate it for each control parameter value to obtain $x_{1}^{*}$, $\ldots$, $x_{N}^{*}$ before running the simulated annealing algorithm to determine a sentinel node set. Therefore, we assessed the runtime of the dynamics as a function of $N$. We remind that we used the implicit Adams method and a fixed total simulation time of $T=15$ (see section IV-B in the main text). We show the runtime for the 20 networks in Fig. S34. The figure indicates that the runtime scales as somewhat more slowly than $O(N^{2})$ as $N$ increases. This result in combination with Fig. S32 supports that the dynamics simulations are not a computational bottleneck as $N$ becomes large when we use our framework for transfer learning assuming no knowledge about the real dynamics.
Figure S34: Runtime of the dynamics on networks. For each pair of dynamics and network, we measured the runtime summed over all the control parameter values. A symbol represents the runtime summed over all 100 control parameter values for a pair of dynamics and network. See the caption of Fig. S32 for the hardware used.